# Cardinality of the sets of dimension functions in ordered structures

### v1

@@ 2. Weak dimension function is a dimension function
Let $\mathcal{M}=(M,\ldots)$ be a structure. A function $\dim:\operatorname{Def}(\mathcal{M})\to\mathbb{N}\cup\{-\infty\}$ on $\mathcal{M}$ is a weak dimension function on $\mathcal{M}$ if conditions (1), $(2)_{1}$ and $(4)_{n}$ for $n>1$ in Definition 1.1 and the following condition (3’), which is equivalent to $(3)_{2}$ in Definition 1.1, is satisfied:
  1. (3’)
@@
Let $\mathcal{M}=(M,\ldots)$ be a structure. Let $\dim:\operatorname{Def}(\mathcal{M})\to\mathbb{N}\cup\{-\infty\}$ be a function. Let $n$ be a positive integer with $n>1$. If $(2)_{1}$, $(2)_{n-1}$ and $(4)_{n}$ in Definition 1.1 holds for $\dim$, then $(2)_{n}$ holds for $\dim$.
@@
Suppose $\dim S_{0}(X)>\dim S_{1}(X)$. We have $\dim(X)=\dim(S_{0}(X))$ by $(4)_{n}$. We want to show
$\dim S_{0}(X\cup Y)=\dim(S_{0}(X)\setminus S_{1}(Y))=\dim X.$
@@
$\leq\dim X$
by $(2)_{n-1}$ and $(4)_{n}$. By $(2)_{n-1}$ and this inequality, we have $\dim(S_{0}(X)\setminus S_{1}(Y))<\dim X$. We have
@@
by $(4)_{n}$ and $(2)_{n-1}$, which is absurd. We have shown $\dim(S_{0}(X)\setminus S_{1}(Y))=\dim X$. This means $\dim S_{0}(X\cup Y)=\dim X$. We also have
@@
Suppose $\dim S_{0}(X)\leq\dim S_{1}(X)$. We have $\dim(X)=\dim(S_{1}(X))+1$ by $(4)_{n}$. We have
@@
by $(2)_{n-1}$. Therefore, we have $\dim S_{0}(X\cup Y)\leq\dim X$. It is obvious $\dim S_{1}(Y)<\dim Y\leq\dim X$. We have
@@ 1. $(5)_{n}$
If $n=2$, property $(3)_{n}$ in Definition 1.1 immediately follows from property (3’) in Definition 2.2.
@@
Suppose $l=n$. Apply property (3’) in Definition 2.2 to $X_{x}^{\Pi_{2}^{n}}\subseteq M^{2}$ for $x\in\Pi_{2}^{n}(X)$. Then, we get $\dim(X_{x}^{\Pi_{2}^{n}})=\dim((X^{\sigma})_{x}^{\Pi_{2}^{n}})$ for every $x\in\Pi_{2}^{n}(X)$. Therefore, we have $X^{\sigma}(2,k)=X(2,k)$ for $0\leq k\leq 2$. Using $(2)_{n}$ and $(5)_{n}$, we can prove $\dim X=\dim X^{\sigma}$ in the same manner as in the previous case. We omit the details. ∎
@@ 3.3. O-minimal case
Let $\mathcal{M}$ be a structure and $\dim:\operatorname{Def}(\mathcal{M})\to\mathbb{N}\cup\{-\infty\}$ be a function satisfying conditions (1), $(2)_{1}$ and $(4)_{n}$ for $n>1$ in Definition 1.1. Let $I$ be a definable subset of $M$ with $\dim I=1$.
@@
Let $\mathcal{M}$, $\dim:\operatorname{Def}(\mathcal{M})\to\mathbb{N}\cup\{-\infty\}$ and $I$ be as in Definition 3.11. The function $\operatorname{Dim}[\dim,I]$ defined in Definition 3.11 satisfies condition $(1)$, $(2)_{1}$ and $(4)_{n}$ for $n>1$ in Definition 1.1.
@@
Satisfaction of (1) and $(2)_{1}$ is obvious. We omit the proof. We prove $(4)_{n}$ by induction on $n$. Note that $(2)_{i}$ holds for $1\leq i<n$ by Lemma 2.3 and the induction hypothesis. We abbreviate $\operatorname{Dim}[\dim,I]$ by $\operatorname{Dim}$ for the simplicity of notations.
Let $\tau\in\{0,1\}^{n}$. First we consider the case in which $X\subseteq I\langle\tau\rangle$. Suppose $\tau(n)=0$. Put $X(i):=S_{i}(X)$ for $i=0,1$. They are obviously definable. Equality ($(4)_{n}$) in $(4)_{n}$ obviously holds by the definition of $\operatorname{Dim}$ for $i=1,2$. Next suppose $\tau(n)=1$. We have $\operatorname{Dim}(X)=\operatorname{Dim}(\Pi(X))$ by the definition of $\operatorname{Dim}$. In this case, we have $X(0)=\Pi(X)$ and $X(1)=\emptyset$. They are definable, and equality ($(4)_{n}$) obviously holds.
Consider arbitrary $X$. Put
@@
Let $\tau\in\{0,1\}^{n}$. By the abuse of notation, let $\Pi(\tau)$ be the element $\tau^{\prime}\in\{0,1\}^{n-1}$ such that $\tau^{\prime}(i)=\tau(i)$ for $1\leq i\leq n-1$. If $\tau(n)=0$, we have
$S_{1}(X\cap\Pi^{-1}(X(1))\cap I\langle\tau\rangle)=S_{1}(X\cap I\langle\tau\rangle)=X(1)\cap I\langle\Pi(\tau)\rangle.$
On the other hand, $S_{0}(X\cap\Pi^{-1}(X(1))\cap I\langle\tau\rangle)=\emptyset$. Hence, we have
$\operatorname{Dim}(X\cap\Pi^{-1}(X(1))\cap I\langle\tau\rangle)=\operatorname{Dim}(X(1)\cap I\langle\Pi(\tau)\rangle)+1$
by the definition of $\operatorname{Dim}$. If $\tau(n)=1$, by $(2)_{n-1}$, we have
$\operatorname{Dim}(X\cap\Pi^{-1}(X(1))\cap I\langle\tau\rangle)=\operatorname{Dim}(\Pi(X\cap I\langle\tau\rangle)\cap X(1))$
$\qquad\leq\operatorname{Dim}(X(1)\cap I\langle\Pi(\tau)\rangle)$
because $\Pi(X\cap I\langle\tau\rangle)\cap X(1)\subseteq X(1)\cap I\langle\Pi(\tau)\rangle$.
By summarizing the above two inequalities, if $\tau_{1},\tau_{2}\in\{0,1\}^{n}$, $\Pi(\tau_{1})=\Pi(\tau_{2})$, $\tau_{1}(n)=0$ and $\tau_{2}(n)=1$, we have
$\operatorname{Dim}(X\cap\Pi^{-1}(X(1))\cap I\langle\tau_{1}\rangle)\geq\operatorname{Dim}(X\cap\Pi^{-1}(X(1))\cap I\langle\tau_{2}\rangle).$
Finally, we obtain
$\operatorname{Dim}(X\cap\Pi^{-1}(X(1)))=\max\{\operatorname{Dim}(X\cap\Pi^{-1}(X(1))\cap I\langle\tau\rangle)\;|\;\tau\in\{0,1\}^{n}\}$
$\qquad=\max\{\operatorname{Dim}(X\cap\Pi^{-1}(X(1))\cap I\langle\tau\rangle)\;|\;\tau\in\{0,1\}^{n},\tau(n)=0\}$
$\qquad=\max\{\operatorname{Dim}(X(1)\cap I\langle\Pi(\tau)\rangle)+1\;|\;\tau\in\{0,1\}^{n},\tau(n)=0\}$
$\qquad=\max\{\operatorname{Dim}(X(1)\cap I\langle\tau^{\prime}\rangle)\;|\;\tau^{\prime}\in\{0,1\}^{n-1}\}+1$
$\qquad=\operatorname{Dim}(X(1))+1.$
We have proved equality ($(4)_{n}$) in $(4)_{n}$ of Definition 1.1 for $i=1$.
Let $\tau\in\{0,1\}^{n}$ with $\tau(n)=0$. We have
$S_{1}(X\cap\Pi^{-1}(X(0))\cap I\langle\tau\rangle)=\emptyset\text{ and }$
$S_{0}(X\cap\Pi^{-1}(X(0))\cap I\langle\tau\rangle)=X(0)\cap\Pi(X\cap I\langle\tau\rangle).$
Therefore, we have
$\operatorname{Dim}(X\cap\Pi^{-1}(X(0))\cap I\langle\tau\rangle)=\operatorname{Dim}(X(0)\cap\Pi(X\cap I\langle\tau\rangle))$
by the definition of $\operatorname{Dim}$. Let $\tau\in\{0,1\}^{n}$ with $\tau(n)=1$. We have
$\operatorname{Dim}(X\cap\Pi^{-1}(X(0))\cap I\langle\tau\rangle)=\operatorname{Dim}(X(0)\cap\Pi(X\cap I\langle\tau\rangle)).$
On the other hand, we get
$\operatorname{Dim}(X(0))$ $=\operatorname{Dim}(X(0)\cap\Pi(X))$
$=\max\{\operatorname{Dim}(X(0)\cap\Pi(X\cap I\langle\tau\rangle))\;|\;\tau\in\{0,1\}^{n}\}$
by $(2)_{n-1}$ because $X=\bigcup_{\tau\in\{0,1\}^{n}}X\cap I\langle\tau\rangle$. Using this equality, we have
$\operatorname{Dim}(X\cap\Pi^{-1}(X(0)))=\max\{\operatorname{Dim}(X\cap\Pi^{-1}(X(0))\cap I\langle\tau\rangle)\;|\;\tau\in\{0,1\}^{n}\}$
$\qquad=\max\{\operatorname{Dim}(X(0)\cap\Pi(X\cap I\langle\tau\rangle))\;|\;\tau\in\{0,1\}^{n}\}$
$\qquad=\operatorname{Dim}(X(0)).$
We have proved equality ($(4)_{n}$) for $i=0$. ∎
@@ 3.4. Weakly o-minimal case
We show that $X(0)$ and $X(1)$ definable. We may reduce to simpler cases by covering $X$ by finitely many simpler definable sets $X_{1},\ldots X_{l}$. In fact, we have $X(1)=\bigcup_{i=1}^{l}X_{i}(1)$ and $X(0)=\Pi(X)\setminus X(1)$ by $(2)_{1}$. If $X_{i}(0)$ and $X_{i}(1)$ are definable for all $1\leq i\leq l$, then $X(0)$ and $X(1)$ are definable. By Claim 3, we may assume that $X$ is a good definable set without loss of generality. Claim 7 implies that $X(0)$ and $X(1)$ are definable. Condition $(4)_{n}$ in Definition 1.1 obviously holds by the definition of $\dim_{\text{w}}$.
We show that $\dim_{\text{w}}$ satisfies condition (3’) in Definition 2.2. Let $X$ be a definable subset of $M^{2}$. By Lemma 2.3, condition $(2)_{2}$ in Definition 1.1 holds. Using $(2)_{2}$ and Claim 3, we may assume that $X$ is a good definable set without loss of generality. We define $\mathcal{I}$, $k_{i}$, $f_{i}(x)$ and $\mathcal{S}_{i}^{k_{i}}(f_{i}(x))$ as in the definition of good definable sets. Since $\mathcal{M}$ is weakly o-minimal, $\Pi(X)$ is a union of a finite set and finitely many definable open convex set. By partitioning $\Pi(X)$, by $(2)_{2}$, we may assume the following:
@@
We have proved that $\dim_{\text{w}}$ satisfies (3’) in Definition 2.2. Proposition 2.5 implies that $\dim_{\text{w}}$ is a dimension function.

### v2

@@ 2. Weak dimension function is a dimension function
Let $\mathcal{M}=(M,\ldots)$ be a structure. A function $\dim:\operatorname{Def}(\mathcal{M})\to\mathbb{N}\cup\{-\infty\}$ on $\mathcal{M}$ is a weak dimension function on $\mathcal{M}$ if conditions (1), $(2)_{1}$ in Definition 1.1 and the following condition (3’), which is equivalent to $(3)_{2}$ in Definition 1.1, and $(4^{\prime})_{n}$ for $n>1$ is satisfied:
  1. $(3^{\prime})$
@@ 2. $(4^{\prime})_{n}$
Let $\mathcal{M}=(M,\ldots)$ be a structure. Let $\dim:\operatorname{Def}(\mathcal{M})\to\mathbb{N}\cup\{-\infty\}$ be a function. Let $n$ be a positive integer with $n>1$. If $(2)_{1}$, $(2)_{n-1}$ in Definition 1.1 and $(4^{\prime})_{n}$ in Definition 2.2 hold for $\dim$, then $(2)_{n}$ holds for $\dim$.
@@
Suppose $\dim S_{0}(X)>\dim S_{1}(X)$. We have $\dim(X)=\dim(S_{0}(X))$ by $(4^{\prime})_{n}$. We want to show
$\dim(S_{0}(X)\setminus S_{1}(Y))=\dim X.$
@@
$=\dim X$
by $(2)_{n-1}$ and $(4^{\prime})_{n}$. By $(2)_{n-1}$ and this inequality, we have $\dim(S_{0}(X)\setminus S_{1}(Y))<\dim X$. We have
@@
by $(4^{\prime})_{n}$ and $(2)_{n-1}$, which is absurd. We have shown $\dim(S_{0}(X)\setminus S_{1}(Y))=\dim X$.
On the other hand, we have
$\dim(S_{0}(Y)\setminus S_{1}(X))\leq\dim S_{0}(Y)\leq\dim Y\leq\dim X$
by $(2)_{n-1}$, $(4^{\prime})_{n}$ and the case hypothesis. We have
$\dim S_{0}(X\cup Y)$ $=\max\{\dim(S_{0}(X)\setminus S_{1}(Y)),\dim(S_{0}(Y)\setminus S_{1}(X))\}=\dim X$
by $(2)_{n-1}$.
We also have
@@
Suppose $\dim S_{0}(X)\leq\dim S_{1}(X)$. We have $\dim(X)=\dim(S_{1}(X))+1$ by $(4^{\prime})_{n}$. We have
@@
by $(2)_{n-1}$ and $(4^{\prime})_{n}$. We get $\dim(S_{0}(Y)\setminus S_{1}(X))\leq\dim Y$ in the same manner. We obtain $\dim(S_{0}(Y)\setminus S_{1}(X))\leq\dim X$ by the case hypothesis. Therefore, we have $\dim S_{0}(X\cup Y)\leq\dim X$. It is obvious $\dim S_{1}(Y)<\dim Y\leq\dim X$ by $(4^{\prime})_{n}$. We have
@@ 1. $(5)_{n}$
If $n=2$, property $(3)_{n}$ in Definition 1.1 immediately follows from property $(3^{\prime})$ in Definition 2.2.
@@
Suppose $l=n$. Apply property $(3^{\prime})$ in Definition 2.2 to $X_{x}^{\Pi_{2}^{n}}\subseteq M^{2}$ for $x\in\Pi_{2}^{n}(X)$. Then, we get $\dim(X_{x}^{\Pi_{2}^{n}})=\dim((X^{\sigma})_{x}^{\Pi_{2}^{n}})$ for every $x\in\Pi_{2}^{n}(X)$. Therefore, we have $X^{\sigma}(2,k)=X(2,k)$ for $0\leq k\leq 2$. Using $(2)_{n}$ and $(5)_{n}$, we can prove $\dim X=\dim X^{\sigma}$ in the same manner as in the previous case. We omit the details. ∎
@@ 3.3. O-minimal case
Let $\mathcal{M}$ be a structure and $\dim:\operatorname{Def}(\mathcal{M})\to\mathbb{N}\cup\{-\infty\}$ be a function satisfying conditions (1), $(2)_{1}$ in Definition 1.1 and $(4^{\prime})_{n}$ in Definition 2.2 for $n>1$ in Definition 1.1. Let $I$ be a definable subset of $M$ with $\dim I=1$.
@@
Let $\mathcal{M}$, $\dim:\operatorname{Def}(\mathcal{M})\to\mathbb{N}\cup\{-\infty\}$ and $I$ be as in Definition 3.11. The function $\operatorname{Dim}[\dim,I]$ defined in Definition 3.11 satisfies condition $(1)$, $(2)_{1}$ and $(4^{\prime})_{n}$ for $n>1$ in Definition 1.1 and Definition 2.2.
@@
Satisfaction of (1) and $(2)_{1}$ is obvious. We omit the proof. We prove $(4^{\prime})_{n}$ by induction on $n$. Note that $(2)_{i}$ holds for $1\leq i<n$ by Lemma 2.3 and the induction hypothesis. We abbreviate $\operatorname{Dim}[\dim,I]$ by $\operatorname{Dim}$ for the simplicity of notations.
Let $X$ be a definable subset of $M^{n}$. Put
@@
We put $X_{\tau}:=X\cap\Pi^{-1}(I\langle\tau\rangle)$ for $\tau\in\{0,1\}^{n-1}$. Define $X_{\tau}(i)$ in the same manner as $X(i)$ for $i=0,1$. It is obvious that $X(i)=\bigcup_{\tau\in\{0,1\}^{n-1}}X_{\tau}(i)$. For every $\tau\in\{0,1\}^{n-1}$, let $\tau[i]\in\{0,1\}^{n}$ such that $\tau[i](j)=\tau(j)$ for $1\leq j<n$ and $\tau[i](n)=i$ for $i=0,1$.
Suppose $X_{\tau}$ enjoys property $(4^{\prime})_{n}$ for every $\tau\in\{0,1\}^{n-1}$. By the definition of $\operatorname{Dim}$ and $(2)_{n-1}$, we have
$\operatorname{Dim}X$ $=\max\{\operatorname{Dim}(X\cap I\langle\tau^{\prime}\rangle)\;|\;\tau^{\prime}\in\{0,1\}^{n}\}$
$=\max\{\max\{\operatorname{Dim}(X\cap I\langle\tau[i]\rangle)\;|\;i=0,1\}\;|\;\tau\in\{0,1\}^{n-1}\}\}$
$=\max\{\operatorname{Dim}X_{\tau}\;|\;\tau\in\{0,1\}^{n-1}\}$
$=\max\{\max\{\operatorname{Dim}X_{\tau}(i)+i\;|\;i=0,1\}\;|\;\tau\in\{0,1\}^{n-1}\}$
$=\max\{\max\{\operatorname{Dim}X_{\tau}(i)\;|\;\tau\in\{0,1\}^{n-1}\}+i\;|\;i=0,1\}$
$=\max\{\operatorname{Dim}X(i)+i\;|\;i=0,1\}.$
Therefore, we may assume that there exists $\tau\in\{0,1\}^{n-1}$ such that $\Pi(X)\subseteq I\langle\tau\rangle$ without loss of generality.
Suppose $\Pi(X)\subseteq I\langle\tau\rangle$ for some $\tau\in\{0,1\}^{n-1}$. Let $X_{i}:=X\cap I\langle\tau[i]\rangle$ for $i=0,1$. Define $X_{0}(i)$ in the same manner as $X(i)$ for $i=0,1$. Put $Y:=\Pi(X_{1})$. By the definition of $\operatorname{Dim}$, we have the following:
$X(1)$ $=X_{0}(1)$
$X(0)$ $=X_{0}(0)\cup(Y\setminus X(1))$
It is obvious from the definition of $\operatorname{Dim}$ that
$\operatorname{Dim}X$ $=\max\{\operatorname{Dim}X_{i}\;|\;i=0,1\}$
$=\max\{\operatorname{Dim}X_{0}(1)+1,\operatorname{Dim}X_{0}(0),\operatorname{Dim}Y\}$
$=\max\{\operatorname{Dim}X(1)+1,\operatorname{Dim}X_{0}(0),\operatorname{Dim}Y\}.$
We consider three separate cases. First suppose $\operatorname{Dim}Y>\operatorname{Dim}X(1)$. We have $\operatorname{Dim}(Y\setminus X(1))=\operatorname{Dim}Y$ by $(2)_{n-1}$. We have $\operatorname{Dim}X(0)=\max\{\operatorname{Dim}X_{0}(0),\operatorname{Dim}Y\}$ by $(2)_{n-1}$. This implies $\operatorname{Dim}X=\max\{\operatorname{Dim}X(0),\operatorname{Dim}X(1)+1\}$.
Next suppose $\operatorname{Dim}Y\leq\operatorname{Dim}X(1)<\operatorname{Dim}X_{0}(0)$. We have $\operatorname{Dim}(Y\setminus X(1))<\operatorname{Dim}Y<\dim X_{0}(0)$. This implies $\operatorname{Dim}X(0)=\operatorname{Dim}X_{0}(0)$. Therefore,
$\max\{\operatorname{Dim}X(0),\operatorname{Dim}X(1)+1\}=\max\{\operatorname{Dim}X_{0}(0),\operatorname{Dim}X(1)+1\}$
$=\max\{\operatorname{Dim}X_{0}(0),\operatorname{Dim}Y,\operatorname{Dim}X(1)+1\}=\operatorname{Dim}X.$
Suppose $\operatorname{Dim}Y\leq\operatorname{Dim}X(1)$ and $\operatorname{Dim}X_{0}(0)\leq\operatorname{Dim}X(1)$. By $(2)_{n-1}$, we get $\operatorname{Dim}X(0)\leq\max\{\dim X_{0}(0),\dim Y\}\leq\dim X(1)$. We get
$\operatorname{Dim}(X)$ $=\max\{\operatorname{Dim}X(1)+1,\operatorname{Dim}X_{0}(0),\operatorname{Dim}(Y)\}$
$=\operatorname{Dim}(X(1))+1=\max\{\operatorname{Dim}X(0),\operatorname{Dim}X(1)+1\}.$
@@ 3.4. Weakly o-minimal case
We show that $X(0)$ and $X(1)$ definable. We may reduce to simpler cases by covering $X$ by finitely many simpler definable sets $X_{1},\ldots X_{l}$. In fact, we have $X(1)=\bigcup_{i=1}^{l}X_{i}(1)$ and $X(0)=\Pi(X)\setminus X(1)$ by $(2)_{1}$. If $X_{i}(0)$ and $X_{i}(1)$ are definable for all $1\leq i\leq l$, then $X(0)$ and $X(1)$ are definable. By Claim 3, we may assume that $X$ is a good definable set without loss of generality. Claim 7 implies that $X(0)$ and $X(1)$ are definable. Condition $(4^{\prime})_{n}$ in Definition 2.2 obviously holds by the definition of $\dim_{\text{w}}$.
We show that $\dim_{\text{w}}$ satisfies condition $(3^{\prime})$ in Definition 2.2. Let $X$ be a definable subset of $M^{2}$. By Lemma 2.3, condition $(2)_{2}$ in Definition 1.1 holds. Using $(2)_{2}$ and Claim 3, we may assume that $X$ is a good definable set without loss of generality. We define $\mathcal{I}$, $k_{i}$, $f_{i}(x)$ and $\mathcal{S}_{i}^{k_{i}}(f_{i}(x))$ as in the definition of good definable sets. Since $\mathcal{M}$ is weakly o-minimal, $\Pi(X)$ is a union of a finite set and finitely many definable open convex set. By partitioning $\Pi(X)$, by $(2)_{2}$, we may assume the following:
@@
We have proved that $\dim_{\text{w}}$ satisfies $(3^{\prime})$ in Definition 2.2. Proposition 2.5 implies that $\dim_{\text{w}}$ is a dimension function.